OpenStudy (anonymous):

Find the slope of the line joining (1/3 , 2/5) and (3/4, 3,8) use fractions not decimals

OpenStudy (anonymous):

no one likes fractions i see

OpenStudy (anonymous):

\[\frac{\frac{3}{8}-\frac{2}{5}}{\frac{3}{4}-\frac{1}{3}}\] is a start

OpenStudy (anonymous):

yes i did that and got 1/40 - 5/12

OpenStudy (anonymous):

\[\frac{-\frac{1}{40}}{\frac{5}{12}}\]

OpenStudy (anonymous):

\[-\frac{1}{40}\times \frac{12}{5}\]

OpenStudy (anonymous):

\[-\frac{3}{50}\]

OpenStudy (anonymous):

the 1/40 is negative?

OpenStudy (anonymous):

definitely

OpenStudy (anonymous):

ok thank you !! :)